OpenStudy (anonymous):

Let f(x) = sin^10(x) − cos(x), 0 <= x<=pi  . Show that there is a solution of f(x) = 0 in (0,pi ) please help

OpenStudy (michele_laino):

the function \(f(x)\) is continuous inside all the real line, furthermore, we have: \(f(0)=-1\), and \(f(\pi)=1\) so what can you conclude?

OpenStudy (anonymous):

I did this, but I didn't how to complete the answer... please, can you just explain to me.. it may come in the final test

OpenStudy (michele_laino):

such \(continuous\) function takes all values inside the set \([-1,1]\), so there exists, inside the interval \([0, \pi]\), at least one point \(x_0\), such that: \(f(x_0)=0\)

OpenStudy (michele_laino):

since such function have to go in a \(continuous\) way, from negative values to positive values

OpenStudy (anonymous):

So, what is the point,that will lead the equation to be zero??? please help

OpenStudy (michele_laino):

the point is that the function is continuous, so, it has to take all values between \(-1<0\) and \(1>0\), hence , such function will be equal to zero at least one time inside \([0,\pi]\)

OpenStudy (anonymous):

thank yooooooooooooooooooooooou so much,,, I got it.......

OpenStudy (michele_laino):

the continuity is the key property

OpenStudy (michele_laino):

:)